OpenStudy (anonymous):

In how many different ways can the top eleven new indie bands be ranked on a top eleven list? The top hit song for each of the eleven bands will compete to receive monetary awards of $1000, $500, $250 and $100, respectively. In how many ways can the awards be given out?

OpenStudy (anonymous):

There are: \[11! = 11*10*9*...*1 = 39916800\] different combinations possible for the top eleven and: \[11!/7! = 11*10*9*8 = 7920\] different award combinations.

OpenStudy (anonymous):

awesome thats what i got after trying a thousand time (smiling).. thought it was wrong. I am getting better i appreciate your help